OpenStudy (anonymous):

Find the solution: 7 cos^-1 x - π = 5 cos^-1 x

OpenStudy (anonymous):

7/(cosx) - Pi = 5/(cosx) >> [7/(cosx)] - [5/(cosx)] = Pi >> 2/(cosx) = Pi >> 2 = cos (Pix) >> cosx = 2/Pi

OpenStudy (anonymous):

If you substitute 2/Pi in the place of cosx you should get the same answer for both sides

OpenStudy (anonymous):

Idk if this helps, but the answer is {0}

OpenStudy (anonymous):

What are you looking for? cosx ?

OpenStudy (anonymous):

Let me write it down, hold on a sec o:

OpenStudy (anonymous):

I'm a bit confused by you "/", are they division signs? o:

OpenStudy (anonymous):

yes, they mean division. To be honest I'm not sure how the answer comes out to be zero, this seemed like the logical answer to me. Sorry :(

OpenStudy (anonymous):

is the -Pi in the denominator ?

OpenStudy (anonymous):

But why were you dividing from the beginning? o:

OpenStudy (chillout):

I guess (s)he though that it was cos^-1 instead of arccos. Well, Going to try to solve it.

OpenStudy (chillout):

By the way, on the left side of the equation is it cos^-1(x-π) or cos^-1(x) - π?

OpenStudy (anonymous):

All she gave me was 7 cos^-1 x -π, so I suppose the second one

hartnn (hartnn):

yes, it must be \(7 cos^{-1} x - π = 5 cos^{-1} x\) Also, \(cos^{-1} x \ne 1/ \cos x\) its inverse cos function, or called 'arccos x' so, lets first try to isolate \(cos^{-1} x\) can you ?

OpenStudy (anonymous):

cox^-1 x=π/2, right?

hartnn (hartnn):

correct! now \(if, cos^{-1} a=b, \implies a=\cos b\) so, \(cos^{-1} x=\pi/2 \implies x = \cos (\pi/2)=..?\)

OpenStudy (anonymous):

x? o:

hartnn (hartnn):

finding solution means finding 'x' so, yes 'x' do you know cos (pi/2) =... ?

OpenStudy (anonymous):

0 OMG THANK YOU.

hartnn (hartnn):

yes, x=0 welcome ^_^